Whats the number in standard form?
#1. 3.41x10^7 #2. 4.05x10^-4
And what is 6,500,000,000 in scientific notation

Found 2 solutions by josgarithmetic, Theo:
Answer by josgarithmetic(39838)

Putting 4.05x10^-4 into standard form, simply perform the multiplication (moving the decimal point correctly).

0.000405


You can determine what to do with the other two examples.

Answer by Theo(13342)

to convert 3.41 * 10^7 to standard form, you are moving the decimal point to the right 7 places.
that gets you 34,100,000.
moving the decimal point 7 places to the right can be viewed as:
moving it 2 places to the right to get 341 * 10^5 and then adding five 0's to get 34,100,000

to convert 4.05 * 10^-4 to standard form, move the decimal point to the left 4 places.
move it to the left 1 place and it becomes .405 * 10^-3
move it to left 3 more places and it becomes .000405

to convert 6,500,000,000 to scientific notation, you want to move the decimal point to the left so that it will be just to the right of the first significant digit.

in order to dot that, you have to move it to the left 9 places (the commas don't count as a space - remove them before doing the move.
you get 6500000000. before moving.
you get 6.5 * 10^9 after moving.

how do you know you did it right?
use your calculator to divide the scientific form by the standard form.
you should get 1.

3.41 * 10^7 / 34100000 = 1, so that was good.
4.05 * 10^-4 / .000405 = 1, so that was good.
6.5 * 10^9 / 6500000000 = 1, so that was good.

if you try to do it the other way (dividing the standard form by the scientific form), you will need to use parentheses.

for example:

34100000 / 3.41 * 10^7 will get you 1 * 10^12.
this is not what what you expected.
but, 34100000 / (3.41 * 10^7) will get you 1.
this is what you expected.